OpenStudy (mendicant_bias):

Alright, my last hurdle to tackle (Systems of Linear Differential Equations), first problem in the section. Posted below momentarily.

OpenStudy (mendicant_bias):

\[\frac{dx}{dt}=-x+y\]

OpenStudy (mendicant_bias):

\[\frac{dy}{dt}=2x\]

OpenStudy (mendicant_bias):

@Kainui , any way you could help me on this? This is my first time dealing with a system of linear diff eqns in any sense. Should I integrate with respect to one or the other, or divide and substitute in? I'm not really sure.

OpenStudy (sidsiddhartha):

ok came back for this last one :)

OpenStudy (sidsiddhartha):

start with diffrentiating the first equation

OpenStudy (sidsiddhartha):

\[\frac{ dx }{ dt }=-x+y\\ \frac{ d^2x }{ dt^2 }=-1+\frac{ dy }{ dt }=2x-1\] getting this?

OpenStudy (mendicant_bias):

back, sorry, was dealing with somebody IRL

OpenStudy (mendicant_bias):

I took a different approach, I took the laplace of both sides of all of them and then substituted in for one and took the inverse laplace and stuff

OpenStudy (mendicant_bias):

But yeah, that makes sense. Goodnight!

OpenStudy (mendicant_bias):

Alright, taking the Laplace of both sides: \[sX(s)-x(0)=-X(s)+Y(s);\]\[sY(s)-y(0)=2X(s)\]

OpenStudy (mendicant_bias):

(Got distracted by stuff, attempting to solve now)

OpenStudy (mendicant_bias):

From (2), \[Y(s)=\frac{2X(s)+y(0)}{s}\]

OpenStudy (mendicant_bias):

Substituting into the first equation, \[sX(s)-x(0)=-X(s)+\frac{2X(s)+y(0)}{s}\]

OpenStudy (mendicant_bias):

Isolating the X(s) term to one side, \[sX(s)+X(s)-\frac{2X(s)}{s}=\frac{y(0)}{s}+x(0)\]

OpenStudy (mendicant_bias):

Now Isolating X(s) entirely,

OpenStudy (mendicant_bias):

\[X(s)\bigg[s+1-\frac{2}{s}\bigg]=\frac{y(0)}{s}+x(0)\]

OpenStudy (mendicant_bias):

Forgot the initial conditions, posting below momentarily

OpenStudy (mendicant_bias):

x(0)=0, y(0)=1

OpenStudy (mendicant_bias):

\[\frac{1}{s\bigg[s+1-\frac{2}{s}\bigg]}\]

OpenStudy (mendicant_bias):

Gonna be a mistake somewhere in there, lol, either way, this doesn't look pretty. Does this look right so far/do you have any suggestions about how to proceed or how I should've proceeded?

OpenStudy (mendicant_bias):

OH

OpenStudy (mendicant_bias):

\[\frac{1}{s^2+s-2}\]

OpenStudy (mendicant_bias):

\[\frac{1}{(s+2)(s-1)}\]

OpenStudy (mendicant_bias):

Alright, know how to move from here.